At t = 0.0s, a runner is moving forward (in the positive direction along the x-axis) at a speed of 7.0 m/s. For 2.0 s the runner accelerates to his left at 0.50 m/s2 and backward at 1.0 m/s2. what is the runners position at t = 2.0 s in terms of the component directions (forward/back and left/right)?
1
Expert's answer
2019-09-25T09:01:34-0400

Way forward "S_1 = 7\cdot2 - \frac{1\cdot2^{2}}{2} = 12 (m),"


path left "S_2 = \frac{0.5\cdot2^{2}}{2} = 1 (m)" ,


therefore, the position of the runner is "forward and left."